OpenStudy (logan13):

Which is equal to 42 • 48? A. 416 B. 410 C. 46 D. 4−6

OpenStudy (anonymous):

What is the dot?

OpenStudy (haseeb96):

just multiply 43 with 48 and get your answer her dot show multiplication sign

OpenStudy (thadyoung):

You can't do this on your calculator? Like, really?

OpenStudy (anonymous):

@thadyoung XDDD

OpenStudy (thadyoung):

If you don't have one: http://www.online-calculator.com/

OpenStudy (anonymous):

But if you multiply those numbers then it doesn't equal any of those answers

OpenStudy (phi):

I assume you mean 4^2 * 4^8 that is 4 times itself 10 times

OpenStudy (anonymous):

The answer to the multiplication isn't one of the options, so it must be something else.

OpenStudy (logan13):

@phi yes

OpenStudy (phi):

do you now that \[ 4^2 = 4\cdot 4 \] ? and \[ 4^8 = 4 \cdot 4\cdot 4\cdot 4\cdot 4\cdot 4\cdot 4\cdot 4\] do you see that putting a little 8 in the upper right is short-hand for multiplying 4 times itself 8 times?

OpenStudy (phi):

so the answer is \[ 4^2 \cdot 4^8 \\ = 4 \cdot 4\cdot 4\cdot 4\cdot 4\cdot 4\cdot 4\cdot 4 \cdot 4 \cdot 4\]

OpenStudy (logan13):

so c

OpenStudy (phi):

I posted the answer. count the number of 4's how many are there being multiplied together?

OpenStudy (logan13):

d